# From a Tiny Slip to a Giant Leap: An LLM-Based Simulation for Fake News Evolution

### v1

@@
Abstract.
With the growing spread of misinformation online, research has increasingly focused on detecting and tracking fake news. However, an overlooked issue is that fake news doesn’t naturally exist in social networks—it often originates from distorted facts or deliberate fabrication by malicious actors. Understanding how true news gradually evolves into fake news is critical for early detection and prevention, reducing its spread and impact. Hence, in this paper, we take the first step toward simulating and revealing this evolution, proposing a Fake News evolUtion Simulation framEwork (FUSE) based on large language models (LLMs). Specifically, we employ LLM as agents to represent individuals in a simulated social network. We define four types of agents commonly observed in daily interactions: spreaders, who propagate information; commentators, who provide opinions and interpretations; verifiers, who check the accuracy of information; and bystanders, who passively observe without engaging. For simulated environment, we model various social network structures, such as high-clustering networks and scale-free networks, to mirror real-world network dynamics. Each day, the agents engage in belief exchanges, reflect on their thought processes, and reintroduce the news accordingly. Given the lack of prior work in this area, we developed a FUSE-EVAL evaluation framework to measure the deviation from true news during the fake news evolution process. The results show that FUSE successfully captures the underlying patterns of how true news transforms into fake news and accurately reproduces previously discovered instances of fake news, aligning closely with human evaluations. Moreover, our work provides insights into the fact that combating fake news should not be delayed until it has fully evolved; instead, prevention in advance is key to achieving better outcomes. We hope our work inspires further research and responsibility in the early detection and prevention of fake news in social networks.  [^1] Our code and data will be released in camera ready version.
Fake News Evolution, LLM-Based Simulation, Agent, Intervention
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems World Wide Webccs: Computing methodologies Artificial intelligence
1. INTRODUCTION
The rapid spread of fake news has become a significant global concern, with social media accelerating the production, consumption, and propagation (Lazer et al., 2018b; Olan et al., 2022). Addressing the evolution of fake news is therefore critical to mitigating its far-reaching effects and is particularly important for fostering a responsible web environment (Aïmeur et al., 2023; Qin, 2020; Adriani, 2019). Most existing studies on fake news focus on detecting misinformation or simulating its spread once it has already been generated (Garimella et al., 2017; Wang et al., 2019). For instance, Piqueira et al. (2020) categorized individuals into four types and used mathematical models to simulate the spread of fake news, as depicted in Figure 1(a), showing the dynamics of different actor populations over time. On a micro-level, Jalili and Perc (2017) defined numerical conditions for opinion change to study fake news dissemination, as shown in Figure 1(b). The above work follows a key assumption common in much of the existing research: fake news inherently exists within social networks.
However, a significantly overlooked issue in these works is that fake news does not naturally exist within social networks. In reality, it may originate from a piece of true news that, over time, becomes distorted, misinterpreted, or altered as it spreads among individuals, eventually evolving into fake news (Guo et al., 2021; Wang et al., 2021; Sunstein, 2014; Herzing and Muck, 2024; Shen et al., 2024) as illustrated in Figure 1(c), or it may be intentionally generated by malicious actors, to grab public attention or serve specific agendas such as political manipulation or financial gain (Kimmel, 2013; Schindler, 2007). Here we define partially evolved fake news as news content that contains elements of factual information but has been manipulated, misrepresented, or selectively edited, resulting in a mixture of both accurate and false details. The dynamic and evolving nature of fake news underscores the need for a deeper understanding of the underlying transformation mechanisms. Addressing this gap is crucial, as understanding these mechanisms can help us design more effective early interventions to prevent fake news from gaining traction within social networks (Schindler, 2007; Shen et al., 2024; DiFonzo and Bordia, 2011).
@@ 1. INTRODUCTION
Figure 1. (a) Macro-level observation of population dynamics based on the mathematical model, categorizing individuals into four types and showing their quantity changes over time. (b) Micro-level visualization of the conventional fake news dissemination model, which assumes fake news inherently exists. (c) Micro-level evolution of fake news, where true news gradually transforms into fake news as it propagates through a network, with content alterations at various stages.
Correspondingly, in this paper, we focus on modeling and understanding fake news formative process, and present the first Fake news evolUtion Simulation framEwork (FUSE). Unlike traditional epidemiological models, which assume predefined transition probabilities between different states of belief (Nekovee et al., 2007; N Zehmakan and Galam, 2020), we employ LLMs to simulate human behaviors, generate sophisticated interactions, reflect on information, and give their own version of news report. For the simulation environment setting, we construct various social network structures, such as high-clustering and scale-free networks, to mirror real-world online interaction patterns. Under this framework, we define four types of agents commonly observed in daily interactions: spreaders, who propagate information; commentators, who provide opinions and interpretations; verifiers, who check the accuracy of information; and bystanders, who passively observe without engaging. Each agent represents an individual with unique attributes, including name, age, personality traits, and education level. Each day, agents interact with its neighbor peers, exchanging beliefs. They adjust their beliefs based on their reasoning, prior experiences, and the views of others. Agents are equipped with hierarchical memory: short-term memory records daily interactions, while long-term memory stores accumulated knowledge and prior beliefs. Additionally, agents use a reflective reasoning process, enabling them to reconsider past experiences and update their believed news content dynamically.
@@
To our best knowledge, there is no prior work on language-based evaluation of fake news evaluation. Hence, we introduce FUSE-EVAL, which tracks information distortion and provides quantitative insights into how true news evolves into fake news. It is a multi-dimensional evaluation framework that analyzes the degree of deviation between news after evolution and the original news, which includes aspects such as Sentiment Shift (SS), New Information Introduced (NII), Certainty Shift (CS), STylistic Shift (STS), Temporal Shift(TS), and Perspective Deviation (PD). Our FUSE framework is also validated through comprehensive simulation experiments, demonstrating a strong alignment with real-world observations from prior research. First, our simulations reveal a clear accumulation distortion effect as news spreads through the network, with information gradually deviating further from its original form (De Fleur, 1962; de Paula et al., 2024; Nian et al., 2022). Additionally, in high-clustering networks, true news evolves into fake news more rapidly compared to scale-free and random networks, consistent with previous studies (Trpevski et al., 2010; Lind et al., 2007). Furthermore, political fake news spreads significantly faster than topics such as terrorism, natural disasters, science, urban legends, or financial information, which also aligns with earlier findings (Rubin, 2022; Lazer et al., 2018a).
@@
To construct a responsible web environment, our research underscores the importance of strategic interventions during the early stages of fake news evolution. Rather than waiting until misinformation has widely spread, we introduce an official agent that actively intervenes when the degree of information deviation reaches a critical threshold. This agent issues timely official statements, incorporating reliable sources to slow down or counteract the evolution of fake news. By providing a trusted narrative early in the process, this approach helps correct circulating misinformation, reducing its impact and limiting further spread. These findings offer practical insights into the role of timely, authoritative interventions for effective misinformation governance.
Our contribution can be summarized in the following ways:
@@
  • We propose FUSE, an LLM-based simulation framework to investigate how true news gradually evolves into fake news, considering the roles of different types of agents and various social network structures observed in real-world.
@@
  • We introduce FUSE-EVAL, a novel evaluation framework to measure the deviation from true information during the news evolution process.
  • We conduct experiments to show that our simulation aligns with real-world observations and previous studies.
@@
  • We propose and evaluate multiple intervention strategies aimed at mitigating the spread of fake news during its evolution.
2. RELATED WORK
2.1. Fake News Evolution
@@
Recent research into fake news evolution has focused on how misinformation spreads and transforms over time. Zhang et al. (2013) found that rumors evolve as they are repeatedly modified, becoming shorter and more shareable, while Guo et al. (2021) empirically tracked fake news evolution, noting how sentiment and text similarity change as truth transitions into misinformation. Xia et al. (2020) developed a sentiment analysis pipeline that addresses sarcasm and out-of-domain input, helping track shifts in public opinion related to fake news. Other studies have emphasized structural and behavioral aspects of fake news propagation. Zhao et al. (2024) proposed a dynamic learning method that captures temporal changes in rumor propagation, revealing how rumor patterns evolve. Wang et al. (2021) demonstrated that slight content changes during the COVID-19 pandemic aided the spread of misinformation, while Li et al. (2016) examined how user behaviors, particularly the role of verified accounts, influence the evolution and dissemination of rumors. These studies underscore the dynamic nature of fake news evolution and the need for comprehensive approaches.
@@
However, there has not been a detailed and comprehensive study on how true news evolves into fake news, with only some superficial linguistic analyses (Zhang et al., 2013; Guo et al., 2021).
2.2. Simulations of Fake News Propagation
Fake news propagation modeling plays a vital role in understanding the dynamics of misinformation spread and is key to developing intervention strategies such as early warning systems (Garimella et al., 2017), misinformation blocking (Song et al., 2015), and truth confrontation (Wang et al., 2019). Current models can be broadly categorized into three types: epidemic models, point process-based models, and diffusion models (Sun et al., 2023). Epidemic models, inspired by the spread of infectious diseases, categorize users as susceptible, infected, or recovered. The SIR model (Zhu and Wang, 2017) is widely used to simulate how misinformation spreads and shows how network structures affect its speed and reach (López et al., 2024). Point process-based models treat fake news spread as random events over time, capturing the temporal dynamics and predicting key influencer engagement (Chen et al., 2019; Gao et al., 2019). Diffusion models focus on individual thresholds, where users decide to spread misinformation based on peer influence, making them effective for studying echo chambers and tight-knit communities (Jalili and Perc, 2017).
Our work extends these models, using LLMs to simulate the evolution of true news into fake news.
2.3. LLM-based Agents for Social Simulation
Agent-based modeling and simulation focuses on representing complex systems by simulating individual agents and their interactions within a dynamic environment (Macal and North, 2005). The integration of LLMs into agent-based modeling has significantly expanded the potential of such simulations, particularly in digital environments where LLM agents excel in natural language tasks (Chen et al., 2023b, a; Liu et al., 2024a). Moreover, LLM-based agents can simulate human-like intelligence, actively engaging in tasks through self-directed planning and decision-making (Xi et al., 2023), and their capabilities extend beyond text generation (Yuhan et al., 2023). These strengths have led to the broad adoption of LLM agents across various domains (Liu et al., 2024b; Park et al., 2022; Zhang et al., 2024; Li et al., 2023; Park et al., 2023; Lin et al., 2023; Kovač et al., 2023; Jinxin et al., 2023; M. Bran et al., 2024; Boiko et al., 2023), making them a new paradigm for simulation by embedding human-level intelligence into agents.
@@ 2.3. LLM-based Agents for Social Simulation
In more specific applications, LLM agents have been employed to simulate social media dynamics. For instance, Törnberg et al. (2023) used LLMs with agent-based modeling to investigate social media algorithms and provide insights into real-world phenomena. Similarly, Park et al. (2022) demonstrated that LLM agents could generate social media content indistinguishable from human-generated material. These developments showcase the growing potential of LLM agents in modeling human social behavior at a collective level. To the best of our knowledge, our work is one of the first to apply LLMs in simulating the evolution of fake news within an agent-based framework.
3. METHODOLOGY
3.1. Problem Formulation
In this study, we aim to simulate the gradual evolution of true news into fake news within a social network by leveraging LLMs as agents. We construct a simulation environment with a pool of $N$ agents $\mathcal{A}=(a_{1},\ldots,a_{N})$, each endowed with a unique persona that includes their role type (spreader, commentator, verifier, or bystander), personality traits, demographic information, and initial opinion toward a news topic.
At time $t=0$, a piece of true news $S_{0}$ is introduced into the network. The agents are connected according to a predefined social network structure, $\mathcal{G}=(\mathcal{A},\mathcal{E})$, which may represent high-clustering, scale-free, or random networks to reflect real-world dynamics. On each day $t=1,2,\ldots,T$, agents interact with their neighbors as defined by the network structure. During these interactions, they exchange information and opinions, influenced by their personas and prior knowledge. After interactions, agents reintroduce the news content considering the new information received. The evolution of the news content held by agent $a_{i}$ at time $t$, denoted as $S_{i}^{t}$, depends on their previous news content and the information received from their neighbors. This update process is defined by:
(1) $S_{i}^{t}=f(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}),$
where $f(\cdot)$ represents the agent’s information processing function, influenced by their persona $\mathcal{P}_{i}$ and the interactions with neighboring agents $\mathcal{N}_{i}$.
@@ 3.1. Problem Formulation
Figure 2. Our FUSE framework simulates news evolution by equipping each agent with role-based decision-making capabilities. Propagation Role-aware agents (PRA) process true news through interactions within the news evolution simulator (NES), where their role identities shape how they engage with the news.
@@
Our objective is to analyze how the true news $S_{0}$ transforms over time due to the agents’ interactions and personal biases. This simulation allows us to research the roles of different agent types, network structures, news topics, and individual personality traits in evolution. By tracking the changes in $S_{i}^{t}$, we can quantify the deviation from the original news and understand the factors influencing the transformation into fake news.
3.2. Our Simulation Framework
@@
As depicted in Figure 2, our FUSE framework integrates two core components: the Propagation Role-Aware agents (PRA) and the News Evolution Simulator (NES). The PRA module equips each agent with role-based decision-making capabilities, enabling them to process and disseminate information according to their assigned roles and personal attributes. The NES constructs the interaction environment, simulating the social network through which news propagates and evolves. Within the PRA module, each agent is powered by an LLM and is assigned a specific role type along with attributes. It influences how agents interpret information, interact with others, and update their opinions. Daily interactions are facilitated by the NES, which determines which agents interact based on the predefined social network structure, $\mathcal{G}=(\mathcal{A},\mathcal{E})$. The NES simulates various network types. During each day, agents engage with their neighbors, exchanging news content and opinions influenced by their roles and personal attributes. In scenarios where interventions are introduced—such as official announcements after the deviation level of news reaches a set threshold, aiming to correct misinformation and release credible information sources to relevant agents. At the end of each day, the simulation progresses by a one-time step, and the agents’ states are updated.
3.3. Propagation Role-Aware Agent
@@
The PRA is the core component of our FUSE framework, designed to simulate individual human behaviors in the context of news evolution. By equipping agents with specific roles and personal attributes, we aim to closely mimic the diversity and complexity of human interactions in social networks.
3.3.1. Personal Information.
@@
3.3.2. Role-Specific Behaviors.

### v2

# The Stepwise Deception: Simulating the Evolution from True News to Fake News with LLM Agents

@@
Abstract
With the growing spread of misinformation online, understanding how true news evolves into fake news has become crucial for early detection and prevention. However, previous research has often assumed fake news inherently exists rather than exploring its gradual formation. To address this gap, we propose FUSE (Fake news evolUtion Simulation framEwork), a novel Large Language Model (LLM)-based simulation approach explicitly focusing on fake news evolution from real news. Our framework model a social network with four distinct types of LLM agents commonly observed in daily interactions: spreaders who propagate information, commentators who provide interpretations, verifiers who fact-check, and bystanders who observe passively to simulate realistic daily interactions that progressively distort true news. To quantify these gradual distortions, we develop FUSE-EVAL, a comprehensive evaluation framework measuring truth deviation along multiple linguistic and semantic dimensions. Results show that FUSE effectively captures fake news evolution patterns and accurately reproduces known fake news, aligning closely with human evaluations. Experiments demonstrate that FUSE accurately reproduces known fake news evolution scenarios, aligns closely with human judgment, and highlights the importance of timely intervention at early stages. Our framework is extensible, enabling future research on broader scenarios of fake news.
The Stepwise Deception: Simulating the Evolution from
True News to Fake News with LLM Agents
Yuhan Liu1, Zirui Song2, Juntian Zhang1, Xiaoqing Zhang1, Xiuying Chen2thanks:   Corresponding authors., Rui Yan1,3,411footnotemark: 1 1Gaoling School of Artificial Intelligence, Renmin University of China, 2MBZUAI, 3Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MoE 4School of Artifcial Intelligence, Wuhan University yuhan.liu@ruc.edu.cn
1 Introduction
@@
Figure 1: (a) Macro-level observation of population dynamics based on the mathematical model, categorizing individuals into four types and showing their quantity changes over time. (b) The micro-level conventional fake news dissemination model assumes that fake news inherently exists. (c) Micro-level evolution of fake news, where true news gradually evolves into fake news during network propagation with content alterations at various stages.
The rapid spread of fake news has become a significant global concern Lazer et al. (2018a); Olan et al. (2022). Prior research predominantly addresses fake news detection or simulates the spread of misinformation after its initial appearance Garimella et al. (2017); Wang et al. (2019). For instance, Piqueira et al. (2020) categorized individuals into four types and used mathematical models to simulate the spread of fake news, as depicted in Figure 1(a). On a micro-level, Jalili and Perc (2017) defined numerical conditions for opinion change to study fake news dissemination, as shown in Figure 1(b). However, these models typically assume fake news as inherently existing entities within social networks, ignoring how misinformation originates or evolves over time.
In contrast, fake news may originate from true news that becomes distorted or misinterpreted over time, eventually evolving into fake news Guo et al. (2021); Shen et al. (2024) as illustrated in Figure 1(c). This evolutionary process is critically underexplored despite its significance for effective early interventions. Recognizing this gap, our work explicitly adopts the definition of fake news from prior influential research Lazer et al. (2018b), focusing specifically on scenarios where factual information incrementally transforms into misinformation during its dissemination (Figure 1(c)). We define this transitional content as partially evolved fake news, characterized by a mix of accurate and distorted elements.
Specifically, we propose the Fake news evolUtion Simulation framEwork (FUSE), the first comprehensive approach employing LLM agents to simulate how real news progressively evolves into fake news within different social network structures (e.g., high-clustering, scale-free, and random networks). The simulation consists of four distinct agent roles commonly found in real-world interactions: spreaders, who disseminate information; commentators, who interpret content; verifiers, who assess factual accuracy; and bystanders, who observe without active participation. Each agent engages daily, exchanging beliefs, reevaluating information, and contributing to incremental content distortions. Our agents incorporate hierarchical memory structures, combining short-term interactions and long-term knowledge, allowing realistic reflective reasoning processes and dynamic content adaptation.
@@
Given the absence of prior work on language-based evaluation of fake news evolution, we introduce FUSE-EVAL, a novel multidimensional evaluation framework that quantifies the deviation of evolved news from its original form across multiple dimensions, including Sentiment Shift (SS), New Information Introduced (NII), Certainty Shift (CS), STylistic Shift (STS), Temporal Shift(TS), and Perspective Deviation (PD). Our comprehensive experiments validate FUSE’s strong alignment with real-world observations from prior research. The results reveal three key findings: (1) news exhibits clear accumulation distortion effects, where content progressively deviates from its original form during spread de Paula et al. (2024); (2) true news evolution to fake news occurs more rapidly in high-clustering networks than in scale-free or random networks Trpevski et al. (2010); (3) political news shows significantly faster evolution rates compared to other topics (terrorism, natural disasters, science, and finance) Lazer et al. (2018b).
@@
To construct a responsible online environment, our research reveals the importance of strategic interventions during the early stages of fake news evolution. Rather than waiting until fake news has widely spread, we introduce an official agent that intervenes when information deviation reaches critical thresholds, issuing authoritative statements with reliable sources to counteract misinformation spread. This early intervention approach demonstrates the effectiveness of timely, authoritative responses in misinformation governance.
Our contributions can be summarized as follows:
@@
$\bullet$ Versatile Framework. We propose FUSE, an LLM-based simulation framework to investigate how true news gradually evolves into fake news, and validate through experiments that our framework successfully reproduces real-world phenomena by considering different types of agents and various social network structures.
@@
$\bullet$ Comprehensive Evaluation. We introduce FUSE-EVAL, a novel multidimensional framework to measure the deviation from true news during news evolution.
@@
$\bullet$ Practical Insights. We propose and evaluate multiple intervention strategies aimed at mitigating the spread of fake news during its evolution.
2 Related Work
Fake News Evolution
@@ 2 Related Work
Recent research into fake news evolution has focused on how misinformation spreads and transforms over time. Zhang et al. (2013) found that rumors evolve as they are repeatedly modified, becoming shorter and more shareable, while Guo et al. (2021) empirically tracked fake news evolution, noting how sentiment and text similarity change as truth transitions into misinformation. Xia et al. (2020) proposed a sentiment analysis pipeline to track public opinion shifts in fake news by detecting sarcasm. Other studies have emphasized structural and behavioral aspects of fake news propagation. Zhao et al. (2024) proposed a dynamic method that captures temporal changes in rumor propagation, revealing how rumor patterns evolve. Wang et al. (2021) demonstrated slight news content changes during the COVID-19 pandemic, while Li et al. (2016) examined how user behaviors, particularly the role of verified accounts, influence the evolution of rumors. FPS Liu et al. (2024) and TED Liu et al. (2025) uses a multi-agent system to study the propagation and detection of fake news.
@@
However, there has not been a detailed and comprehensive study on how true news evolves into fake news, with only some superficial linguistic analyses Zhang et al. (2013); Guo et al. (2021).
LLMs as Agents
@@
Agent-based modeling simulates complex systems through individual agents’ interactions in dynamic environments Macal and North (2005). The integration of LLMs has enhanced these simulations by enabling natural language processing capabilities Zhang et al. (2025); Chen et al. (2023b, a) and human-like intelligence in planning and decision-making Xi et al. (2023). This has led to widespread adoption across various domains Li et al. (2023); Park et al. (2023); Liu et al. (2025); Jin et al. (2025), establishing LLM agents as a new paradigm for human-level intelligence simulation. In more specific applications, LLM agents have been employed to simulate social media dynamics. For instance, Törnberg et al. (2023) used them to investigate social media algorithms and provide insights into real-world phenomena, while Park et al. (2022) demonstrated their ability to generate human-like social media content. Our work extends this approach by being one of the first to apply LLM agents in simulating fake news evolution.
3 Methodology
@@
Figure 2: Our FUSE framework simulates news evolution by equipping each agent with role-based decision-making capabilities. Propagation Role-aware agents (PRA) process true news through interactions within the news evolution simulator (NES), where their role identities shape how they engage with the news.
3.1 Problem Formulation
We simulate the gradual evolution of true news into fake news using LLMs as agents within a social network, consistent with the definition of fake news provided by prior research Lazer et al. (2018b); Guo et al. (2021). The simulation consists of $N$ agents $\mathcal{A}=(a_{1},\ldots,a_{N})$, each endowed with a unique persona defining their Behavior role, personality traits, and demographic information.
At time $t=0$, true news $S_{0}$ is introduced into the network. The agents are connected according to a predefined social network structure $\mathcal{G}=(\mathcal{A},\mathcal{E})$, which may represent high-clustering, scale-free, or random networks to reflect real-world dynamics. On each day $t=1,2,\ldots,T$, agents interact with their neighbors $\mathcal{N}_{i}$, exchanging information and opinions based on their personas and prior knowledge. After interactions, agents process and reintroduce the news content based on their updated beliefs. The evolution of the news content for agent $a_{i}$ with a personal profile $\mathcal{P}_{i}$ at time $t$, denoted as $S_{i}^{t}$, is defined by:
$S_{i}^{t}=f(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}),$ (1)
where $f(\cdot)$ represents the agent’s information processing function.
@@ 3.1 Problem Formulation
Through this simulation, we analyze how the true news $S_{0}$ transforms over time due to agents’ interactions and personal biases, examining the impact of agent types, network structures, and individual traits on the evolution of fake news.
3.2 Our Simulation Framework
@@
As depicted in Figure 2, our FUSE framework consists of two core components: the Propagation Role-Aware agents (PRA) and the News Evolution Simulator (NES). The PRA module empowers agents with role-based decision-making capabilities, while the NES establishes the interaction environment, simulating the social network through which news propagates and evolves. Within the PRA module, each agent is powered by an LLM and characterized by a specific role type and personal attributes, which govern their information processing, interaction patterns, and opinion updates. The NES facilitates daily interactions through a predefined social network structure, $\mathcal{G}=(\mathcal{A},\mathcal{E})$, simulating various network types to reflect different social dynamics.
During each simulation day, agents engage with their network neighbors, exchanging news content and opinions shaped by their roles and attributes. When news content deviates beyond a set threshold, intervention mechanisms—such as official announcements—are triggered to provide credible information and correct potential misinformation. The simulation advances daily with updated agent states, tracking the evolution of news content through the network.
3.3 Propagation Role-Aware Agent
@@
The PRA is designed to simulate individual human behaviors in news evolution by equipping agents with specific roles and personal attributes, aiming to mirror the diversity and complexity of human interactions in social networks.
3.3.1 Personal Information
@@
3.3.2 Role-Specific Behaviors